"It is claimed that a car is driven on the average less than 28700 kilometers per year with a population standard deviation of 2100 kilometers. To test this claim, a random sample of 58 car owners are asked to keep a record of the kilometers they travelled. Would you agree with this claim if the random sample showed an average of 22500 kilometers? Use a 0.01 level of significance.






What are the given? Write only the number. :




population mean:Blank 1 km




population standard deviation:Blank 2 km




sample size:Blank 3




sample mean:Blank 4 km




level of significance:Blank 5






What is the critical value?




z:Blank 6






What is the value of the calculated z? Round your answer to the nearest hundredths.




z:Blank 7"

1
Expert's answer
2022-05-17T14:48:16-0400

population mean: "28700" km

population standard deviation: "2100" km

sample size:"58"

sample mean: "22500" km

level of significance: "0.01"

What is the critical value?

"z_c=-2.3263"

What is the value of the calculated z?


"z=\dfrac{\bar{x}-\mu}{\sigma/\sqrt{n}}=\dfrac{22500-28700}{2100/\sqrt{58}}\approx-22.48"

The following null and alternative hypotheses need to be tested:

"H_0:\mu\ge28700"

"H_1:\mu<28700"

This corresponds to a left-tailed test, for which a z-test for one mean, with known population standard deviation will be used.

Since it is observed that "z=-22.48<-2.3263= z_c," it is then concluded that the null hypothesis is rejected.

Using the P-value approach: The p-value is "p=P(z<-22.48)\approx0," and since "p=0<0.01=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu"

is less than 28700, at the "\alpha = 0.01" significance level.